Assume all sales are one-time credit sales with a probability of collection of 96 percent. The variable cost per unit is $1.67, the sales price per unit is $4.99, and the monthly interest rate is 1.4 percent. What is the NPV of a credit sale of one extra unit?

A) $3.073

B) $3.026

C) $2.981

D) $2.936

E) $2.892
Answer #1

Expected revenue = 4.99 * 96%

Expected revenue = $4.7904

Expected profit = 4.7904 - 1.67 = $3.1204

NPV = Expected profit/(1 + r)

NPV = 3.1204/(1 + 0.014)

NPV = $3.0773175542

Option A is correct: $3.073

Can you please upvote? Thank You :-)